# Removal and Selection: Improving RGB-Infrared Object Detection via Coarse-to-Fine Fusion

### v3

@@ Abstract
Object detection in visible (RGB) and infrared (IR) images has been widely applied in recent years. Leveraging the complementary characteristics of RGB and IR images, the object detector provides reliable and robust object localization from day to night. Existing fusion strategies directly inject RGB and IR images into convolution neural networks, leading to inferior detection performance. Since the RGB and IR features have modality-specific noise, these strategies will worsen the fused features along with the propagation. Inspired by the mechanism of human brain processing multimodal information, this work introduces a new coarse-to-fine perspective to purify and fuse two modality features. Specifically, following this perspective, we design a Redundant Spectrum Removal module to coarsely remove interfering information within each modality and a Dynamic Feature Selection module to finely select the desired features for feature fusion. To verify the effectiveness of the coarse-to-fine fusion strategy, we construct a new object detector called Removal and Selection Detector (RSDet). Extensive experiments on three RGB-IR object detection datasets verify the superior performance of our method.
[…]
1 Introduction
With the development of various sensors, multimodal images are collected and applied in more and more application fields LI et al. (2021); Cao et al. (2015). Among them, visible (RGB) and infrared (IR) sensors are widely utilized due to their complementary imaging characteristics. Specifically, visible images can reflect rich texture and color information of the object but are greatly affected by low-light environments. In contrast, infrared images can reflect the thermal information of the object and can clearly reflect the outline of the object under poor illumination conditions, but lack the texture details of the object. In recent years, researchers have focused on the visible and infrared feature-level fusion  Yuan et al. (2022); Sun et al. (2022); Yuan and Wei (2023), which can help achieve better performance on downstream tasks (e.g., object detection).
@@ 1 Introduction
In RGB-IR object detection, the effective feature fusion method of RGB and IR images is crucial. Most existing RGB-IR object detection methods extract the modality-specific features from RGB and IR images independently, and then directly perform addition or concatenation operations on these features Liu et al. (2016a); Li et al. (2018); Cao et al. (2019), as shown in Figure 1(a). Without explicit cross-modal fusion, the “Late fusion” strategy is therefore limited in learning the complementary information, resulting in inferior performance. To further explore the optimal fusion strategies, many researchers have considered “Halfway fusion” strategy to design an interaction module between different modality features. For instance, Zhou et al. Zhou et al. (2020) construct the DMAF module to tap the difference between RGB and IR modalities which brings more useful information at the feature level. Xie et al. Xie et al. (2022) introduce a novel dynamic cross-modal module that aggregates local and global information from RGB and infrared modalities. Although these methods have achieved encouraging improvements, they ignore the negative impact of the modality-noise along with the propagation, which would fail to achieve complementary fusion.
Actually, when confronted with multi-modal information, our brains first establish rules to eliminate interfering information, and then meticulously select the desired information, which has been modeled in a cognition theory (“Attenuation Theory” Treisman (1964)). Such a way can be summarized as a coarse-to-fine process, which motivates us to introduce a new perspective to fuse RGB and IR features. As shown in Figure 1(c), we design a new fusion strategy called “Coarse-to-Fine Fusion” to achieve complementary feature fusion. “Coarse” indicates that our method begins with the filter out the interfering information and thus can coarsely remove the irrelevant spectrum. To this end, since the redundant information in an image also exists in its frequency spectrum, we propose a Redundant Spectrum Removal (RSR) module to filter out coarsely in the frequency domain. Specifically, we convert each source image into frequency space and introduce a dynamic filter to adaptively reduce irrelevant spectrum within RGB and IR modalities. As for “Fine”, it indicates that our fusion process conducts finely select features after the coarse removal. Therefore, we design a Dynamic Feature Selection (DFS) module to meticulously select the desired features between RGB and IR modalities. In this module, we gate different scale features required for object detection by exploring a mixture of scale-aware experts. Figure 2 gives an example of Coarse-to-Fine fusion strategy. To evaluate the effectiveness of the coarse-to-fine strategy, we build a novel framework for RGB-IR object detection called Removal and Selection Detector (RSDet).
@@
  • We present a new coarse-to-fine perspective to fuse RGB and IR features. Inspired by the mechanism of human brain processing multimodal information, we coarsely remove the interfering information and finely select desired features for fusion.
@@
  • Following the coarse-to-fine fusion strategy, we propose a Redundant Spectrum Removal module, which introduces a dynamic spectrum filter to adaptively reduce irrelevant information in the frequency domain. Besides, we also design a Dynamic Feature Selection module, which utilize a mixture of scale-aware experts to gate different scale features for final fusion.
@@
  • To verify the effectiveness of coarse-to-fine strategy, we build a novel object detector based on it for RGB-IR object detection. Extensive experiments on three public RGB-IR object detection datasets demonstrate that our proposed method achieves state-of-the-art performance.
2 Related Works
2.1 Multispectral Object Detection
@@
In recent years, thanks to the development of deep learning technology and several visible and infrared datasets being proposed Hwang et al. (2015); Zhang et al. (2020); Jia et al. (2021), the multispectral object detection task has gradually attracted attention. To fully explore the effective information between visible and infrared images, some researchers focus on the complementarity between the two modalities starting from the illumination conditions. Guan et al. Guan et al. (2019) and Li et al. Li et al. (2019) first propose the illumination-aware modules to allow the object detectors to adjust the fusion weight based on the predicted illumination conditions. Moreover, Zhou  et al. Zhou et al. (2020) analyze and address the modality imbalance problems by designing two feature fusion modules called DMAF and IAFA. Recently, an MSR memory module Kim et al. (2022) was introduced to enhance the visual representation of the single modality features by recalling the multispectral modality features, which enables the detector to encode more discriminative features. Yuan et al. Yuan et al. (2024) propose a transformer-based RGB-IR object detector to further improve the object detector performance. Different from these works, we present a new coarse-to-fine fusion strategy to fuse RGB and IR features for multispectral object detection task.
@@
2.2 Shared-Specific Representation Learning
Shared and specific representation learning is first explored in the Domain Separation Network Bousmalis et al. (2016) for unsupervised domain adaptation. It uses a shared-weight encoder to capture shared features and a private encoder to capture domain–specific features. Sanchez et al. Sanchez et al. (2020) explored further shared and specific feature disentanglement representation, and found it is useful to perform downstream tasks such as image classification and retrieval. Recently, van et al. van Amsterdam et al. (2023) improved the performance of action segmentation by disentangling the latent features into shared and modality-specific components. Furthermore, Wang et al. Wang et al. (2023) proposed the ShaSpec model handled missing modalities problems. Shared-specific representation learning has shown great performance and effectiveness in feature learning. However, few RGB-IR object detection models explicitly exploit shared-specific representation. In this paper, we introduce shared-specific representation learning between RGB and IR modality features to implement our coarse-to-fine fusion strategy.
2.3 Mixture of Experts
The Mixture-of-Experts (MoE) model Jacobs et al. (1991); Jordan and Jacobs (1994) has demonstrated the ability to dynamically adapt its structure based on varying input conditions. Several studies have been dedicated to the theoretical exploration of MoE Lepikhin et al. (2021); Nie et al. (2022); Zhou et al. (2022), focusing on enhancing model quality, efficient training and router mechanisms. Besides, some researchers also concentrated on leveraging the MoE model for specific tasks. For example, Gross et al. Gross et al. (2017) observed that a hard mixture-of-experts model can be efficiently trained to good effect on large scale multi-label prediction tasks. Cao et al. Cao et al. (2023a) proposed a mixture of local-to-global experts (MoE-Fusion) mechanism by integrating MoE structure into image fusion tasks. Chen et al. Chen et al. (2023) addressed the multi-task learning challenge by implementing a cooperative and specialized mechanism among experts. In our proposed method, the MoE model is integrated into the multispectral object detection task. Specifically, we establish multi-scale experts to gate modality features at different scales and leverage its dynamic fusion mechanism to facilitate complementary fusion.
3 Method
3.1 Coarse-to-Fine Fusion
The design of the “Coarse-to-Fine Fusion” strategy is shown in Figure 3, “Coarse” corresponding to the Redundant Spectrum Removal (RSR) module to filter out coarsely in the frequency domain and ‘Fine’ corresponding to the Dynamic Feature Selection (DFS) module to meticulously select the desired features between RGB-IR modalities. Since the two modality features are usually intersection, we introduce disentangled representation learning Sanchez et al. (2020) to purify and decouple them for complementary fusion. Thus, we integrate above two modules into the shared-specific structures to implement our coarse-to-fine fusion. Specifically, we first perform the RSR module to remove interfering information. Then, we introduce the shared-specific structure to extract the two modality-specific multiscale features $C_{\text{I-spe}}$ and $C_{\text{V-spe}}$, which uses ResNet50 He et al. (2016) as the backbone network. As for the shared features $C_{\text{sha}}$, we also employ several Resblocks as the feature extractor. After that, these different scale features $C_{\text{I-spe}}$ and $C_{\text{V-spe}}$ are dynamically aggregated by exploring a mixture of scale-aware experts and obtain the final specific feature $C_{\text{spe}}$ by DFS module. Finally, the specific feature $C_{\text{spe}}$ and the shared feature $C_{\text{sha}}$ are added together for the final object detection task, which can be expressed as :
$C=C_{\text{sha}}+C_{\text{spe}}.$ (1)
In the next sections, we will further introduce the proposed RSR module and DFS module.
3.2 Redundant Spectrum Removal
For the “Coarse Removal”, since the frequency domain obeys the nature of global modeling, we propose a Redundant Spectrum Removal (RSR) module to filter out coarsely in the frequency domain. Specifically, we convert each source image into frequency space and predict a dynamic filter to adaptively reduce irrelevant spectrum within RGB and IR modalities. Specifically, the paired RGB image $V\in\mathbb{R}^{H\times W\times 3}$ and IR image $I\in\mathbb{R}^{H\times W}$ are token as the RSR input. They are first subjected to the $\operatorname{DFT}(\cdot)$ operation and get the frequency domain image $F_{I}(u,v)$ and $F_{V}(u,v)$:
@@ 3.2 Redundant Spectrum Removal
Then, we perform the $\operatorname{top}K$ operation on $M_{m}$ to get the binary filters $H_{I}(u,v)$ and $H_{V}(u,v)$. In specific, the positions corresponding to the $\operatorname{top}K$ values are set to 1 and the others to 0:
@@
Finally, we apply $H(u,v)$ to filter the frequency domain image $F(u,v)$, like $F_{I}^{{}^{\prime}}(u,v)=F_{I}(u,v)\otimes H_{I}(u,v)$ and $F_{V}^{{}^{\prime}}(u,v)=F_{V}(u,v)\otimes H_{V}(u,v)$ and subject the Inverse Discrete Fourier Transform $\operatorname{IDFT}(\cdot)$ to transform image back to the spatial domain. Thus, we obtain the final irrelevant spectrum removed images $I^{{}^{\prime}}$ and $V^{{}^{\prime}}$.
@@
3.3 Dynamic Feature Selection
@@
Table 1: Detection results ($\text{MR}^{\text{-2}}$, in%) under ‘All-dataset’ settings of different pedestrian distances, occlusion levels, and light conditions (Day and Night) on the KAIST dataset. The pedestrian distances consist of ‘Near’ (115 $\leq$ height), ‘Medium’ (45 $\leq$ height $<$ 115) and ‘Far’ (1 $\leq$ height $<$ 45), while occlusion levels consist of ‘None’ (never occluded), ‘Partial’ (occluded to some extent up to half) and ‘Heavy’ (mostly occluded). IoU $=0.5$ is set for evaluation. The best results are highlighted in red and the second-place are highlighted in blue.
Methods Backbone Near Medium Far None Partial Heavy Day Night All
ACF Hwang et al. (2015) VGG16 28.74 53.67 88.20 62.94 81.40 88.08 64.31 75.06 67.74
Halfway Fusion Liu et al. (2016a) VGG16 8.13 30.34 75.70 43.13 65.21 74.36 47.58 52.35 49.18
FusionRPN+BF Konig et al. (2017) VGG16 0.04 30.87 88.86 47.45 56.10 72.20 52.33 51.09 51.70
IAF R-CNN Li et al. (2019) VGG16 0.96 25.54 77.84 40.17 48.40 69.76 42.46 47.70 44.23
IATDNN+IASS Guan et al. (2019) VGG16 0.04 28.55 83.42 45.43 46.25 64.57 49.02 49.37 48.96
CIAN Zhang et al. (2019a) VGG16 3.71 19.04 55.82 30.31 41.57 62.48 36.02 32.38 35.53
MSDS-R-CNN Li et al. (2018) VGG16 1.29 16.19 63.73 29.86 38.71 63.37 32.06 38.83 34.15
AR-CNN Zhang et al. (2019b) VGG16 0.00 16.08 69.00 31.40 38.63 55.73 34.36 36.12 34.95
MBNet Zhou et al. (2020) ResNet50 0.00 16.07 55.99 27.74 35.43 59.14 32.37 30.95 31.87
TSFADet Yuan et al. (2022) ResNet50 0.00 15.99 50.71 25.63 37.29 65.67 31.76 27.44 30.74
CMPD Li et al. (2022) ResNet50 0.00 12.99 51.22 24.04 33.88 59.37 28.30 30.56 28.98
CAGTDet Yuan et al. (2024) ResNet50 0.00 14.00 49.40 24.48 33.20 59.35 28.79 27.73 28.96
3.4 Removal and Selection Detector (RSDet)
To evaluate the effectiveness of our Coarse-to-Fine fusion strategy, we embed it into an existing object detection framework. In specific, we utilize a two-stage detector Faster R-CNN Ren et al. (2017) as our base model and replace its backbone with our strategy to construct a new object detector called RSDet. Other modules such as Region Proposal Network (RPN) and R-CNN head remain unchanged.

### v4

@@ Abstract
Object detection in visible (RGB) and infrared (IR) images has been widely applied in recent years. Leveraging the complementary characteristics of RGB and IR images, the object detector provides reliable and robust object localization from day to night. Most existing fusion strategies directly input RGB and IR images into deep neural networks, leading to inferior detection performance. However, the RGB and IR features have modality-specific noise, these strategies will exacerbate the fused features along with the propagation. Inspired by the mechanism of the human brain processing multimodal information, in this paper, we introduce a new coarse-to-fine perspective to purify and fuse two modality features. Specifically, following this perspective, we design a Redundant Spectrum Removal module to coarsely remove interfering information within each modality and a Dynamic Feature Selection module to finely select the desired features for feature fusion. To verify the effectiveness of the coarse-to-fine fusion strategy, we construct a new object detector called the Removal and Selection Detector (RSDet). Extensive experiments on three RGB-IR object detection datasets verify the superior performance of our method. The source code and results are available at https://github.com/Zhao-Tian-yi/RSDet.git
Index Terms:
RGB-Infrared Object Detection, Coarse-to-Fine Fusion Strategy, Multisensory Fusion, Mixture of Experts.
I Introduction
Object detection on single modality is one of the fundamental tasks in computer vision, garnering significant attention and finding applications across diverse fields such as surveillance[1, 2], remote sensing[3, 4, 5], and autonomous driving[6, 7]. However, object detection methods reliant solely on one single modality have shown susceptibility to various challenges [8], like illumination, weather, adversarial attack, occlusion, etc. Therefore, the multiple sensor modalities fusion method has come into view[9, 10, 11, 12, 13]. With the development of various sensors, multimodal images are collected and applied in more and more application fields. Among them, visible (RGB) and infrared (IR) sensors are widely utilized due to their complementary imaging characteristics. Specifically, visible images can reflect rich texture and color information of the object but are greatly affected by low-light environments. In contrast, infrared images can reflect the thermal information of the object and can clearly reflect the outline of the object under poor illumination conditions, but lack the texture details of the object. In recent years, researchers have focused on the visible and infrared feature-level fusion  [14, 15, 16, 11], which can help achieve better performance on downstream tasks (e.g., object detection).
@@ I Introduction
In RGB-IR object detection, the effective feature fusion method of RGB and IR images is crucial. Most existing RGB-IR object detection methods extract the modality-specific features from RGB and IR images independently, and then directly perform addition or concatenation operations on these features [17, 18, 19], as shown in Figure 1(a). Without explicit cross-modal fusion, the “Late fusion” strategy is therefore limited in learning the complementary information, resulting in inferior performance. To further explore the optimal fusion strategies, many researchers have considered the “Halfway fusion” strategy to design an interaction module between different modality features [20, 21, 22], as shown in Figure 1(b). For instance, Zhou et al. [20] construct the DMAF module to tap the difference between RGB and IR modalities which brings more useful information at the feature level. Xie et al. [21] introduce a novel dynamic cross-modal module that aggregates local and global information from RGB and IR modalities, etc. Although these methods have achieved encouraging improvements, they ignore the negative impact of the modality-noise along with the propagation, which would fail to achieve complementary fusion.
Actually, when confronted with multi-modal information, our brains initially establish rules to filter out interfering information and then meticulously select the desired information, a process that has been modeled in cognitive theory (“Attenuation Theory” [23]). This approach can be likened to a coarse-to-fine process, inspiring us to introduce a fresh perspective for fusing RGB and IR features. As shown in Figure 1(c), we design a new fusion strategy called “Coarse-to-Fine Fusion” to achieve complementary feature fusion.
“Coarse” indicates that our method begins with the filter out the interfering information and thus can coarsely remove the irrelevant spectrum. To this end, since the redundant information in an image also exists in its frequency spectrum [24], we propose a Redundant Spectrum Removal (RSR) module to filter out coarsely in the frequency domain. Specifically, we convert each source image into frequency space and introduce a dynamic filter to adaptively reduce irrelevant spectrum within RGB and IR modalities. As for “Fine”, it indicates that our fusion strategy conducts finely select features after the coarse removal. We design a Dynamic Feature Selection (DFS) module to meticulously select the desired features between RGB and IR modalities. In this module, we weighted different scale features required for object detection by exploring a mixture of scale-aware experts. Figure 2 gives an example of our Coarse-to-Fine fusion strategy. To evaluate the effectiveness of the coarse-to-fine strategy, we build a novel framework for RGB-IR object detection called Removal and Selection Detector (RSDet).
@@
  • We propose a new coarse-to-fine perspective to fuse RGB and IR features. Inspired by the mechanism of the human brain processing multimodal information, we coarsely remove the interfering information and finely select desired features for fusion.
@@
  • Following the coarse-to-fine fusion strategy, we propose a Redundant Spectrum Removal module, which introduces a dynamic spectrum filter to adaptively reduce irrelevant information in the frequency domain.
  • We also design a Dynamic Feature Selection module, which utilizes a mixture of scale-aware experts to weigh different scale features for the RGB-IR feature fusion.
@@
  • To verify the effectiveness of the coarse-to-fine fusion strategy, we build a novel framework for RGB-IR object detection. Extensive experiments on three public RGB-IR object detection datasets demonstrate our proposed method achieves state-of-the-art performance.
II Related Works
II-A RGB-IR Object Detection
@@
In recent years, thanks to the development of deep learning technology and several visible and infrared datasets being proposed [25, 26, 27], the RGB-IR object detection (also known as multispectral object detection) task has gradually attracted more and more attention. To fully explore the effective information between visible and infrared images, some researchers focus on the complementarity between the two modalities starting from the illumination conditions. Guan et al. [28] and Li et al. [29] first propose the illumination-aware modules to allow the object detectors to adjust the fusion weight based on the predicted illumination conditions. Moreover, Zhou  et al. [20] analyze and address the modality imbalance problems by designing two feature fusion modules called DMAF and IAFA. Recently, an MSR memory module [30] was introduced to enhance the visual representation of the single modality features by recalling the RGB-IR modality features, which enables the detector to encode more discriminative features. Yuan et al. [31] propose a transformer-based RGB-IR object detector to further improve the object detector performance. Different from these works, We are inspired by the classic theory of cognitive psychology, Attenuation Theory, to emulate the way the human brain processes information from multiple sources. We propose a coarse-to-fine fusion strategy to fuse RGB and IR features for the RGB-IR object detection task.
II-B Shared-Specific Representation Learning
Shared and specific representation learning is first explored in the Domain Separation Network [32] for unsupervised domain adaptation. It uses a shared-weight encoder to capture shared features and a private encoder to capture domain-specific features. Sanchez et al. [33] explored further shared and specific feature disentanglement representation, and found it is useful to perform downstream tasks such as image classification and retrieval. Recently, van et al. [34] improved the performance of action segmentation by disentangling the latent features into shared and modality-specific components. Furthermore, Wang et al. [35] proposed the ShaSpec model handled missing modalities problems. Shared-specific representation learning has shown great performance and effectiveness in feature learning. However, few RGB-IR object detection models explicitly exploit shared-specific representation. In this paper, we introduce shared-specific representation learning between RGB and IR modality features to implement our coarse-to-fine fusion strategy.
II-C Mixture of Experts
The Mixture-of-Experts (MoE) model [36, 37] has demonstrated the ability to dynamically adapt its structure based on varying input conditions. Several studies have been dedicated to the theoretical exploration of MoE [38, 39, 40, 41], focusing on the sparsity, training effectiveness, router mechanisms, enhancing model quality, etc. Besides, some researchers also concentrated on leveraging the MoE model for specific downstream tasks. For example, Gross et al. [42] observed that a hard mixture-of-experts model can be efficiently trained to good effect on large-scale multilabel prediction tasks. Cao et al. [43] proposed a mixture of local-to-global experts (MoE-Fusion) mechanisms by integrating MoE structure into image fusion tasks. Chen et al. [44] addressed the multi-task learning challenge by implementing a cooperative and specialized mechanism among experts. In our proposed method, the MoE model is integrated into the RGB-IR object detection task. Specifically, we establish multi-scale experts to gate modality features at different scales and leverage its dynamic fusion mechanism to facilitate complementary fusion.
III Method
III-A Coarse-to-Fine Fusion
The inspiration for the proposed “Coarse-to-Fine” Fusion strategy stems from the cognitive models of human processing of multiple sources of information, named Selective Attention Theory in cognitive psychology (e.g., “Broadbent’s filter model” [45] and “reisman’s Attenuation Model” [23]), which is one of the cornerstones of attention mechanism theory in cognitive psychology. The Selective Attention Theory posits that when the human brain processes multiple external stimuli, it first filters out or attenuates some unimportant or even useless information. Then, it further processes the information in a more refined manner before storing it in working memory, as illustrated in Figure 3. The Attenuating Filter corresponds to the Redundant Spectrum removal module (RSR), which effectively filters out unimportant or interfering information based on some properties and rules when confronted with multi-sensory information. The Hierarchy of Analyzers corresponds to our proposed Dynamic Feature module (DFS), engaging in meticulous, higher-level analysis and processing to extract valuable insights. Finally, the processed information enters the brain’s working memory, corresponding to the output of the fused feature by our “Coarse-to-Fine Fusion” strategy and fed into the detection network.
According to the above inspiration, our “Coarse-to-Fine” Fusion strategy is designed as shown in Figure 4, “Coarse” corresponding to the Redundant Spectrum Removal (RSR) module to filter out coarsely in the frequency domain and ‘Fine’ corresponding to the Dynamic Feature Selection (DFS) module to meticulously select the desired features between RGB-IR modalities. Since the two modality features are usually intersection, we introduce disentangled representation learning [33] to purify and decouple them for complementary fusion. Thus, we integrate the above two modules into the shared-specific structures to implement our coarse-to-fine fusion. Specifically, we first perform the RSR module to remove interfering information. Then, we introduce the shared-specific structure to extract the two modality-specific multiscale features $C_{\text{I-spe}}$ and $C_{\text{V-spe}}$, which uses ResNet50 [46] as the backbone network. As for the shared features $C_{\text{sha}}$, we also employ several Resblocks as the feature extractor. After that, these different scale features $C_{\text{I-spe}}$ and $C_{\text{V-spe}}$ are dynamically aggregated by the proposed mixture of scale-aware experts and obtain the final specific feature $C_{\text{spe}}$ in the DFS module. Finally, the specific feature $C_{\text{spe}}$ and the shared feature $C_{\text{sha}}$ are added together for the final object detection task, which can be expressed as:
$C=C_{\text{sha}}+C_{\text{spe}}.$ (1)
Figure 3: Illustration of Treisman’s Attenuation Model
@@ III-A Coarse-to-Fine Fusion
III-B Redundant Spectrum Removal
For the “Coarse Removal”, since the frequency domain obeys the nature of global modeling, we propose a Redundant Spectrum Removal (RSR) module to filter out coarsely in the frequency domain. we first transform each source image into the frequency domain and predict a dynamic filter to adaptively reduce irrelevant spectrum within RGB and IR modalities. Specifically, the paired RGB image $V\in\mathbb{R}^{H\times W\times 3}$ and IR image $I\in\mathbb{R}^{H\times W\times 1}$ are token as the RSR module input. They are first subjected to the Discrete Fourier Transform ($\operatorname{DFT}(\cdot)$) and get the frequency domain image $F_{I}(u,v)$ and $F_{V}(u,v)$:
@@ III-B Redundant Spectrum Removal
Then, for adaptively reducing irrelevant spectrum according to different inputs by the amplitude information of the original image, we utilize the $\operatorname{top}K$ operation on $M_{m}$. In specific, the positions don’t correspond to the $\operatorname{top}K$ values are set to 0:
@@
we apply $H(u,v)$ to filter the frequency domain image $F(u,v)$ for reducing irrelevant spectrum:
$F_{I}^{{}^{\prime}}(u,v)=F_{I}(u,v)\otimes H_{I}(u,v)$ (5)
$F_{V}^{{}^{\prime}}(u,v)=F_{V}(u,v)\otimes H_{V}(u,v).$
Finally, the filtered $F_{I}^{{}^{\prime}}(u,v)$ and $F_{V}^{{}^{\prime}}(u,v)$ are subjected to the Inverse Discrete Fourier Transform $\operatorname{IDFT}(\cdot)$ to transform the image back to the spatial domain. Thus, we obtain the final irrelevant spectrum removed images $I^{{}^{\prime}}$ and $V^{{}^{\prime}}$.
@@
III-C Dynamic Feature Selection
@@
III-D Removal and Selection Detector (RSDet)
To evaluate the effectiveness of our Coarse-to-Fine fusion strategy, we embed it into an existing object detection framework. In specific, we utilize a two-stage detector Faster R-CNN [47] as our base model and replace its backbone with our strategy to construct a new object detector called RSDet. Other modules such as Region Proposal Network (RPN) and RoI head remain unchanged.